# Finite energy well-posedness for nonlinear Schrödinger equations with non-vanishing conditions at infinity

### v3

@@ 3.1. Local well-posedness
Let the nonlinearity $f$ be such that Assumption 1.1 is satisfied, $T>0$, the pair $(q_{1}^{\prime},r_{1}^{\prime})$ as in (3.1) and $\psi\in L^{\infty}([0,T];\mathbb{E}({\mathbf{R}}^{2}))$, then the following hold
@@
$\|\mathcal{N}_{2,\infty}(\psi)\|_{L_{t}^{1}L_{x}^{2}}\leq CT\mathcal{L}^{2}\left(\operatorname{supp}(1-\eta(|\psi|)\right)^{\frac{1}{2}}\leq CTZ_{T}(\psi).$
@@
$\|\mathcal{N}_{2,q}(\psi)\|_{L_{t}^{1}L_{x}^{2}}\leq C\||\psi|^{2\alpha}|\psi|(1-\chi(\psi))\|_{L_{t}^{1}L_{x}^{2}}\leq CTZ_{T}(\psi)+CT\|\psi_{q}\|_{L_{t}^{\infty}L_{x}^{2(1+2\alpha)}}^{1+2\alpha}\\ \leq CT\left(Z_{T}(\psi)+Z_{T}(\psi)^{1+2\alpha}\right),$
where $\psi_{q}$ is defined in (2.2), with $\chi$ given in (2.1). In the second last inequality, we used that
(3.6) $|\psi|^{2\alpha+1}(1-\chi(\psi))\leq C\left(\mathbf{1}_{\{0<1-\chi(\psi)\leq 1/4\}}+\left|\psi_{q}\right|^{2\alpha+1}\right),$
@@
[…]
To control $\nabla\mathcal{N}(\psi)$, we observe that by using (2.34) and decomposing $\psi=\psi_{\infty}+\psi_{q}$, see (2.2), it follows
$\left\|\nabla\mathcal{N}(\psi)\right\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\leq CT\|\nabla\psi\|_{L_{t}^{\infty}L_{x}^{2}}+\||\psi_{q}|^{2\alpha}\nabla\psi\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\\ \leq C\left(T+T^{\frac{1}{q_{1}^{\prime}}}Z_{T}(\psi)^{2\alpha}\right)\|\nabla\psi\|_{L_{t}^{\infty}L_{x}^{2}}.$
@@
Exploiting that $\mathbb{E}({\mathbf{R}}^{2})+H^{1}({\mathbf{R}}^{2})\subset\mathbb{E}({\mathbf{R}}^{2})$ from Lemma 2.1, we proceed as before to infer that for a.e. $t\in[0,T]$ it holds
@@
which is justified as identity in $\mathbb{E}({\mathbf{R}}^{3})$ in virtue of the properties of the free solutions from Proposition 2.12 and the fact the non-homogeneous terms is bounded in $L_{t}^{\infty}H_{x}^{1}$ by means of the Strichartz estimates (2.25) and Lemma 3.1.
We anticipate that the continuous dependence on the initial data differs significantly from the classical approach as consequence of the low regularity of the nonlinearity $\mathcal{N}$ combined with the lack of integrability of $\psi$. The constructed solutions are such that $\psi(t)-\psi_{0}\in H^{1}({\mathbf{R}}^{2})$ for all $t$ and hence (3.5) suffices to show local existence. Note that in order to show the continuous dependence on the initial data (3.5) is not sufficient as in general different initial data possesses different far-field behavior, namely belongs to different connected components of $\mathbb{E}$, see also Remark 2.4. Lemma 3.3 upgrades (3.5) to the respective inequality for general initial data.
The following Proposition is stated for positive existence times, the analogous statement for negative times follows by exploiting the time reversal symmetry of (1.1).
@@
Let $d=2$ and $f$ be such that Assumption 1.1 is satisfied. Then,
@@
    for any $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{2})$, there exists $T=T(\mathcal{E}(\psi_{0}))>0$ and a unique strong solution $\psi\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ to (1.1) with $\psi(0)=\psi_{0}$. In particular, $\psi-\psi_{0}\in C([0,T];H^{1}({\mathbf{R}}^{2}))$;
@@
Local existence. We note that $\psi\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ is a strong solution to (1.1) with initial data $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{2})$ iff
@@
for all $t\in[0,T]$. To show existence of a solution $\psi$ it suffices to implement a fixed-point argument for the solution map
@@
Indeed, $\psi(t)=\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u(t)$ satisfies $\psi\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ if $u\in X_{T}$ and $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{2})$. It follows from Proposition 2.12 that $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ and Lemma 2.1 yields that $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$. If $u$ is a fixed-point of (4.2) then $\psi=\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u$ is a local strong solution of (1.1).
Let $\psi_{0}\in\mathbb{E}$ and $R>0$ such that $\mathcal{E}(\psi_{0})\leq R$ and given $M>0$ and $T>0$, we consider the solution map (3.8) defined on the function space
@@
It is straightforward to verify that the space $(X_{T},d_{X})$ is a complete metric space. If $\mathcal{E}(\psi_{0})\leq R$ and $u\in X_{T}$, then thanks to the Minkowski inequality and (2.23) we obtain
@@
To bound $\nabla\Phi(u)$, we apply the Strichartz estimates (2.25) concatenated with (3.4) to obtain
@@
and summing up (3.10), (3.11), we obtain that
@@
We fix $M=\sqrt{2R}$ and notice that there exists $0<T\leq 1$ sufficiently small such that
@@
Hence, we deduce that there exists $T_{1}>0$ such that $\psi_{1}=\psi_{2}$ a.e. on $[0,T_{1}]\times{\mathbf{R}}^{2}$. As $T_{1}$ only depends on $Z_{T}(\psi_{1})$, $Z_{T}(\psi_{2})$, one may iterate the argument to obtain uniqueness of the solution on the interval $[0,T]$.
@@
Let $T^{\ast}(\psi_{0})<+\infty$ and assume that there exist $R>0$ and a sequence $\{t_{n}\}_{n\in{\mathbf{N}}}$ such that $t_{n}\rightarrow T^{\ast}(\psi_{0})$ and $\mathcal{E}(\psi(t_{n}))\leq R$ for all $n\in{\mathbf{N}}$. Then, there exists $n$ sufficiently large such that the local existence statement allows us to uniquely extend the solution to $[0,t_{n}+T(R)]$ with $t_{n}+T(R)>T^{\ast}(\psi_{0})$. This violates the maximality assumption and we conclude that
@@
First, we notice that it follows from the first inequality of (2.36) and the decomposition provided by Lemma 2.1 that
@@
where we used that $||\psi_{2}|-1|\eta(|\psi_{2}|)\in L^{\infty}([0,T];L^{\infty}({\mathbf{R}}^{2})\cap L^{2}({\mathbf{R}}^{2}))$. Indeed, let $\Omega\subset{\mathbf{R}}^{2}$ of finite Lebesgue measure and $f\in L^{\infty}(\Omega)+L^{p}(\Omega)$, then
@@
$\|\mathcal{N}_{2,\infty}(\psi_{1})-\mathcal{N}_{2,\infty}(\psi_{2})\|_{L_{t}^{1}L_{x}^{2}}\leq CT(1+{Z_{T}(\psi_{1})}+{Z_{T}(\psi_{2})})\left\|\psi_{1}-\psi_{2}\right\|_{L_{t}^{\infty}(L_{x}^{\infty}+L_{x}^{2})}.$
Third, arguing as in the proof of Lemma 3.1 and exploiting that $\mathcal{L}^{2}(\operatorname{supp}(\mathcal{N}_{2}(\psi_{i})))\leq\mathcal{E}(\psi_{i})$ we obtain
$\|\mathcal{N}_{2,q}(\psi_{1})-\mathcal{N}_{2,q}(\psi_{2})\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\leq\left\|\mathbf{1}_{\operatorname{supp}(1-\chi(\psi_{1}))\cup\operatorname{supp}(1-\chi(\psi_{2}))}|\psi_{1}-\psi_{2}|\right\|_{L_{t}^{1}L_{x}^{2}}\\ +\left\|\left(|\psi_{1,q}|^{2\alpha}+|\psi_{2,q}|^{2\alpha}\right)|\psi_{1}-\psi_{2}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\\ \leq C(T+T^{\frac{1}{q_{1}^{\prime}}})\left({Z_{T}(\psi_{1})}+{Z_{T}(\psi_{2})}+Z_{T}(\psi_{1})^{2\alpha}+Z_{T}(\psi_{2})^{2\alpha}\right)\|\psi_{1}-\psi_{2}\|_{L_{t}^{\infty}(L_{x}^{\infty}+L_{x}^{2})}.$
Concatenating the Strichartz estimates (2.25) and Lemma 3.3 gives the following.
@@
To prove continuous dependence of the solution, we proceed in the following four steps that compensate for the lack of local Lipschitz regularity of $\nabla\mathcal{N}$ that in general does not hold under Assumptions 1.1.
@@
    By iterating (3.18), we prove that all $0<T<T^{\ast}(\mathcal{O}_{\delta})$ and $\varepsilon>0$, there exists $\delta>0$ such that $d_{\mathbb{E}}(\psi_{0}^{1},\psi_{0}^{\ast})<\delta$ yields
@@
Step 1 We show (3.16). Let us consider the first term on the left hand side of (3.16), by using (3.15) and from Lemma 3.4, we know there exists $\theta>0$ such that
@@
Given $\chi$ defined in (2.1), we define $\chi_{6}(z):=\chi(6z)$. Arguing as in the proof of Lemma 2.6 we notice that
@@
where we used (3.3) in the last inequality. Combining the previous inequalities, we infer that there exists $\theta_{1}>0$
@@
For $T_{1}>0$ sufficiently small, only depending on $M$, inequality (3.16) follows.
@@
$\left|\nabla\mathcal{N}(\psi)\right|\leq C(1+|\psi|^{2\alpha})|\nabla\psi|\leq C(1+|\psi_{q}|^{2\alpha})|\nabla\psi|.$
We apply estimate (2.25) to the non-homogeneous term, where $(q_{1},r_{1}))=(\frac{2(\alpha+1)}{\alpha},2(\alpha+1))$, see also (3.1). We decompose $\nabla\mathcal{N}(\psi_{1})-\nabla\mathcal{N}(\psi^{\ast})$ by means of the functions $G_{\infty},G_{q}$ defined in (2.38) leading to
@@
$\leq\left\|(G_{\infty}+G_{q})(\psi_{1})\left|\nabla\psi_{1}-\nabla\psi^{\ast}\right|\right\|_{N^{0}}$
$\qquad+\left\|\left((G_{\infty}+G_{q})(\psi_{1})-(G_{\infty}+G_{q})(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2}))}$
$\leq\|\nabla\psi_{1}-\nabla\psi^{\ast}\|_{L_{t}^{1}L_{x}^{2}}+\||\psi_{1,q}|^{2\alpha}\left|\nabla\psi_{1}-\nabla\psi^{\ast}\right|\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}$
$+\left\|\left(G_{\infty}(\psi_{1})-G_{\infty}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{q}(\psi_{1})-G_{q}(\psi^{\ast})\right)|\nabla\psi_{1}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}$
@@
$+\left\|\left(G_{\infty}(\psi_{1})-G_{\infty}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{q}(\psi_{1})-G_{q}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}$
@@
$\|\nabla\psi_{1}-\nabla\psi^{\ast}\|_{L^{\infty}([0,T_{2}],L^{2}({\mathbf{R}}^{2}))}\leq d_{\mathbb{E}}(\psi_{0}^{1},\psi_{0}^{\ast})\\ +\left\|\left(G_{\infty}(\psi_{1})-G_{\infty}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{q}(\psi_{1})-G_{q}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}.$
@@
(3.27) $\left\|\left(G_{\infty}(\psi^{\ast})-G_{\infty}(\psi_{n})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{q}(\psi^{\ast})-G_{q}(\psi_{n})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\geq\varepsilon,$
where $\psi_{n}\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ denotes the unique maximal solution with $\psi_{n}(0)=\psi_{0}^{n}$. Inequality (3.16) implies that, up to extracting a subsequence, not relabeled, $\psi_{n}$ converges to $\psi^{\ast}$ a.e. on $[0,T_{1}]\times{\mathbf{R}}^{2}$. If $0<T_{1}<T_{2}$, then set $T_{2}:=T_{1}$. By virtue of Assumption 1.1 on $f$, it follows that $G_{\infty},G_{q}$ are continuous and thus
[…]
@@
$\left\|G_{q}(\psi_{n})\right\|_{L_{t}^{\infty}L_{x}^{q_{1}}({\mathbf{R}}^{2})}$ $\leq C\left\|(1+|\psi_{q,n}|^{2\alpha})(1-\chi(\psi_{n})\right\|_{L_{t}^{\infty}L_{x}^{q_{1}}({\mathbf{R}}^{2})}$

### v4

@@ 3.1. Local well-posedness
Let the nonlinearity $f$ be as in Assumption 1.1 is satisfied, $T>0$, the pair $(q_{1}^{\prime},r_{1}^{\prime})$ as in (3.1) and $\psi\in L^{\infty}([0,T];\mathbb{E}({\mathbf{R}}^{2}))$, then the following hold
@@
$\|\mathcal{N}_{2,\mathrm{bd}}(\psi)\|_{L_{t}^{1}L_{x}^{2}}\leq CT\mathcal{L}^{2}\left(\operatorname{supp}(1-\eta(|\psi|)\right)^{\frac{1}{2}}\leq CTZ_{T}(\psi).$
@@
$\|\mathcal{N}_{2,\mathrm{int}}(\psi)\|_{L_{t}^{1}L_{x}^{2}}\leq C\||\psi|^{2\alpha}|\psi|(1-\chi(\psi))\|_{L_{t}^{1}L_{x}^{2}}\leq CTZ_{T}(\psi)+CT\|\psi_{\mathrm{int}}\|_{L_{t}^{\infty}L_{x}^{2(1+2\alpha)}}^{1+2\alpha}\\ \leq CT\left(Z_{T}(\psi)+Z_{T}(\psi)^{1+2\alpha}\right),$
where $\psi_{\mathrm{int}}$ is defined in (2.2), with $\chi$ given in (2.1). In the second inequality, we used that
(3.6) $|\psi|^{2\alpha+1}(1-\chi(\psi))\leq C\left(\mathbf{1}_{\{0<1-\chi(\psi)\leq 1/4\}}+\left|\psi_{\mathrm{int}}\right|^{2\alpha+1}\right),$
@@
$\mathcal{L}^{2}\left(\left\{x\in\operatorname{supp}(1-\chi(\psi))\,:\,0<1-\chi(\psi)\leq 1/4\right\}\right)\leq{Z_{T}(\psi)^{2}}.$
To control $\nabla\mathcal{N}(\psi)$, we observe that by using (2.33) and decomposing $\psi=\psi_{\mathrm{bd}}+\psi_{\mathrm{int}}$, see (2.2), it follows that
$\left\|\nabla\mathcal{N}(\psi)\right\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\leq CT\|\nabla\psi\|_{L_{t}^{\infty}L_{x}^{2}}+\||\psi_{\mathrm{int}}|^{2\alpha}\nabla\psi\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\\ \leq C\left(T+T^{\frac{1}{q_{1}^{\prime}}}Z_{T}(\psi)^{2\alpha}\right)\|\nabla\psi\|_{L_{t}^{\infty}L_{x}^{2}}.$
@@
Exploiting that $\mathbb{E}({\mathbf{R}}^{2})+H^{1}({\mathbf{R}}^{2})\subset\mathbb{E}({\mathbf{R}}^{2})$ from Lemma 2.1, we proceed as before to infer that for a.e. $t\in[0,T]$ we have
@@
which is justified as identity in $\mathbb{E}({\mathbf{R}}^{3})$ in virtue of the properties of the free solutions from Proposition 2.12 and the fact the non-homogeneous terms is bounded in $L_{t}^{\infty}H_{x}^{1}$ by means of the Strichartz estimate (2.24) and Lemma 3.1.
We anticipate that the continuous dependence on the initial data will differ significantly from the classical approach as consequence of the low regularity of the nonlinearity $\mathcal{N}$ combined with the lack of integrability of $\psi$. The constructed solutions are such that $\psi(t)-\psi_{0}\in H^{1}({\mathbf{R}}^{2})$ for all $t$ and hence (3.5) suffices to show local existence. Note that in order to show the continuous dependence on the initial data (3.5) is not sufficient as in general different initial data possess different far-field behavior, namely belongs to different connected components of $\mathbb{E}$, see also Remark 2.4. Lemma 3.3 upgrades (3.5) to the respective inequality for general initial data.
The following Proposition is stated for positive existence times, the analogous statement for negative times follows from the time reversal symmetry of (1.1).
@@
Let $d=2$ and $f$ be as in Assumption 1.1.
@@
    For any $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{2})$, there exist $T=T(\mathcal{E}(\psi_{0}))>0$ and a unique strong solution $\psi\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ to (1.1) with $\psi(0)=\psi_{0}$. In particular, we have that $\psi-\psi_{0}\in C([0,T];H^{1}({\mathbf{R}}^{2}))$;
@@
Local existence. We note that $\psi\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ is a strong solution to (1.1) with initial data $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{2})$ if and only if
@@
for all $t\in[0,T]$. To show the existence of a solution $\psi$ it suffices to implement a fixed-point argument for the solution map
@@
Specifically, let $\psi_{0}\in\mathbb{E}$ and $R>0$ such that $\mathcal{E}(\psi_{0})\leq R$ and given $M>0$ and $T>0$, we consider the solution map (3.8) defined on the function space
@@
It is straightforward to verify that the space $(X_{T},d_{X})$ is a complete metric space.
Note that $\psi(t)=\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u(t)$ satisfies $\psi\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ for $u\in X_{T}$ and $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{2})$. Indeed, it follows from Proposition 2.12 that $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ and Lemma 2.1 yields that $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$. If $u$ is a fixed-point of (4.2) then $\psi=\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u$ is a local strong solution of (1.1).
If $\mathcal{E}(\psi_{0})\leq R$ and $u\in X_{T}$, then thanks to the Minkowski inequality and (2.22) we obtain
@@
To bound $\nabla\Phi(u)$, we apply the Strichartz estimate (2.24) concatenated with (3.4) to obtain
@@
and summing up (3.10) and (3.11), we obtain that
@@
We fix $M=\sqrt{2R}$ and notice that there exists a sufficiently small $0<T\leq 1$ small such that
@@
Hence, we deduce that there exists $T_{1}>0$ such that $\psi_{1}=\psi_{2}$ a.e. on $0,T_{1}]\times{\mathbf{R}}^{2}$. As $T_{1}$ only depends on $Z_{T}(\psi_{1})$, $Z_{T}(\psi_{2})$, one may iterate the argument to obtain uniqueness of the solution on the interval $[0,T]$.
@@
Let $T^{\ast}(\psi_{0})<+\infty$ and assume that there exist $R>0$ and a sequence $\{t_{n}\}_{n\in{\mathbf{N}}}$ such that $t_{n}\rightarrow T^{\ast}(\psi_{0})$ and $\mathcal{E}(\psi(t_{n}))\leq R$ for all $n\in{\mathbf{N}}$. Then, there exists a sufficiently large $n$ such that the local existence statement allows us to uniquely extend the solution to $0,t_{n}+T(R)]$ with $t_{n}+T(R)>T^{\ast}(\psi_{0})$. This violates the maximality assumption and we conclude that
@@
First, we notice that from the first inequality of (2.35) and the decomposition provided by Lemma 2.1, it follows
@@
where we used that $||\psi_{2}|-1|\eta(|\psi_{2}|)\in L^{\infty}([0,T];L^{\infty}({\mathbf{R}}^{2})\cap L^{2}({\mathbf{R}}^{2}))$. Indeed, let the set $\Omega\subset{\mathbf{R}}^{2}$ of finite Lebesgue measure and $f\in L^{\infty}(\Omega)+L^{p}(\Omega)$, then
@@
$\|\mathcal{N}_{2,\mathrm{bd}}(\psi_{1})-\mathcal{N}_{2,\mathrm{bd}}(\psi_{2})\|_{L_{t}^{1}L_{x}^{2}}\leq CT(1+{Z_{T}(\psi_{1})}+{Z_{T}(\psi_{2})})\left\|\psi_{1}-\psi_{2}\right\|_{L_{t}^{\infty}(L_{x}^{\infty}+L_{x}^{2})}.$
Third, arguing as in the proof of Lemma 3.1 and using $\mathcal{L}^{2}(\operatorname{supp}(\mathcal{N}_{2}(\psi_{i})))\leq\mathcal{E}(\psi_{i})$ we obtain
$\|\mathcal{N}_{2,\mathrm{int}}(\psi_{1})-\mathcal{N}_{2,\mathrm{int}}(\psi_{2})\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\leq\left\|\mathbf{1}_{\operatorname{supp}(1-\chi(\psi_{1}))\cup\operatorname{supp}(1-\chi(\psi_{2}))}|\psi_{1}-\psi_{2}|\right\|_{L_{t}^{1}L_{x}^{2}}\\ +\left\|\left(|\psi_{1,\mathrm{int}}|^{2\alpha}+|\psi_{2,\mathrm{int}}|^{2\alpha}\right)|\psi_{1}-\psi_{2}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\\ \leq C(T+T^{\frac{1}{q_{1}^{\prime}}})\left({Z_{T}(\psi_{1})}+{Z_{T}(\psi_{2})}+Z_{T}(\psi_{1})^{2\alpha}+Z_{T}(\psi_{2})^{2\alpha}\right)\|\psi_{1}-\psi_{2}\|_{L_{t}^{\infty}(L_{x}^{\infty}+L_{x}^{2})}.$
Concatenating the Strichartz estimate (2.24) and Lemma 3.3 gives the following.
@@
To prove continuous dependence of the solution, we proceed in the following four steps that compensate for the lack of local Lipschitz regularity of $\nabla\mathcal{N}$ that in general does not hold under Assumption 1.1.
@@
    By iterating (3.18), we prove that for all $0<T<T^{\ast}(\mathcal{O}_{\delta})$ and $\varepsilon>0$, there exists $\delta>0$ such that $d_{\mathbb{E}}(\psi_{0}^{1},\psi_{0}^{\ast})<\delta$ yields
@@
Step 1 We show (3.16). Let us consider the first term on the left hand side of (3.16). By using (3.15) and from Lemma 3.4, we know there exists $\theta>0$ such that
@@
Given $\chi$ satisfying (2.1), we define $\chi_{6}(z):=\chi(6z)$. Arguing as in the proof of Lemma 2.6 we notice that
@@
where we used (3.3) in the last inequality. Combining the previous inequalities, we infer that there exists $\theta_{1}>0$ such that
@@
For $T_{1}>0$ sufficiently small, depending only on $M$, inequality (3.16) follows.
@@
$\left|\nabla\mathcal{N}(\psi)\right|\leq C(1+|\psi|^{2\alpha})|\nabla\psi|\leq C(1+|\psi_{\mathrm{int}}|^{2\alpha})|\nabla\psi|.$
We apply estimate (2.24) to the non-homogeneous term, where $(q_{1},r_{1}))=(\frac{2(\alpha+1)}{\alpha},2(\alpha+1))$, see also (3.1). We decompose $\nabla\mathcal{N}(\psi_{1})-\nabla\mathcal{N}(\psi^{\ast})$ by means of the functions $G_{\mathrm{bd}},G_{\mathrm{int}}$ defined in (2.37) leading to
@@
$\leq\left\|(G_{\mathrm{bd}}+G_{\mathrm{int}})(\psi_{1})\left|\nabla\psi_{1}-\nabla\psi^{\ast}\right|\right\|_{N^{0}}$
$\qquad+\left\|\left((G_{\mathrm{bd}}+G_{\mathrm{int}})(\psi_{1})-(G_{\mathrm{bd}}+G_{\mathrm{int}})(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2}))}$
$\leq\|\nabla\psi_{1}-\nabla\psi^{\ast}\|_{L_{t}^{1}L_{x}^{2}}+\||\psi_{1,\mathrm{int}}|^{2\alpha}\left|\nabla\psi_{1}-\nabla\psi^{\ast}\right|\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}$
$+\left\|\left(G_{\mathrm{bd}}(\psi_{1})-G_{\mathrm{bd}}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{\mathrm{int}}(\psi_{1})-G_{\mathrm{int}}(\psi^{\ast})\right)|\nabla\psi_{1}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}$
@@
$+\left\|\left(G_{\mathrm{bd}}(\psi_{1})-G_{\mathrm{bd}}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{\mathrm{int}}(\psi_{1})-G_{\mathrm{int}}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}$
@@
$\|\nabla\psi_{1}-\nabla\psi^{\ast}\|_{L^{\infty}([0,T_{2}],L^{2}({\mathbf{R}}^{2}))}\leq d_{\mathbb{E}}(\psi_{0}^{1},\psi_{0}^{\ast})\\ +\left\|\left(G_{\mathrm{bd}}(\psi_{1})-G_{\mathrm{bd}}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{\mathrm{int}}(\psi_{1})-G_{\mathrm{int}}(\psi^{\ast})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}.$
@@
(3.27) $\left\|\left(G_{\mathrm{bd}}(\psi^{\ast})-G_{\mathrm{bd}}(\psi_{n})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{1}L_{x}^{2}}+\left\|\left(G_{\mathrm{int}}(\psi^{\ast})-G_{\mathrm{int}}(\psi_{n})\right)|\nabla\psi^{\ast}|\right\|_{L_{t}^{q_{1}^{\prime}}L_{x}^{r_{1}^{\prime}}}\geq\varepsilon,$
where $\psi_{n}\in C([0,T];\mathbb{E}({\mathbf{R}}^{2}))$ denotes the unique maximal solution with $\psi_{n}(0)=\psi_{0}^{n}$. Inequality (3.16) implies that, up to extracting a subsequence, not relabeled, $\psi_{n}$ converges to $\psi^{\ast}$ a.e. on $0,T_{1}]\times{\mathbf{R}}^{2}$. If $0<T_{1}<T_{2}$, then set $T_{2}:=T_{1}$. By virtue of Assumption 1.1 on $f$, it follows that $G_{\mathrm{bd}},G_{\mathrm{int}}$ are continuous and thus
$\left|\left(G_{\mathrm{bd}}(\psi^{\ast})-G_{\mathrm{bd}}(\psi_{n})\right)\right|\,|\nabla\psi^{\ast}|$ $\rightarrow 0\quad\text{a.e. in}\quad[0,T_{2}]\times{\mathbf{R}}^{2},$
$|G_{\mathrm{int}}(\psi^{\ast})-G_{\mathrm{int}}(\psi_{n})|\,|\nabla\psi^{\ast}|$ $\rightarrow 0\quad\text{a.e. in}\quad[0,T_{2}]\times{\mathbf{R}}^{2}.$
@@
$\left\|G_{\mathrm{int}}(\psi_{n})\right\|_{L_{t}^{\infty}L_{x}^{q_{1}}({\mathbf{R}}^{2})}$ $\leq C\left\|(1+|\psi_{q,n}|^{2\alpha})(1-\chi(\psi_{n})\right\|_{L_{t}^{\infty}L_{x}^{q_{1}}({\mathbf{R}}^{2})}$